OpenStudy (anonymous):

what is the simplified form of the expression [m^-1 m^5/m^-2]^3

OpenStudy (anonymous):

Do you know: \[\large a^x \cdot a^y = a^{x + y}\]

OpenStudy (anonymous):

In numerator, you have -1 and 5 in the powers of m.. When base is same(here base is m), then exponents add when you multiply...

OpenStudy (anonymous):

Just add the powers in the numerator part...

OpenStudy (anonymous):

These types of problems lose me... I dont understand why they put letters with numbers. im dyslexic so it makes it 100xs harder.. so bare with me.

OpenStudy (anonymous):

I am here with you, don't worry.. Clearly tell me, where you are not getting, I will start from that point?

OpenStudy (anonymous):

the negative numbers do i add them together or do i subtract the pos. from neg. or am i over thinking completely?

OpenStudy (anonymous):

See, if you are adding or subtracting, there is no difference.. But don't interchange the sign anywhere.. You have -1 and 5 as powers.. Now you want to add them, so you will do like this : (-1) + (5) Just see the addition sign in between them...

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so 6?

OpenStudy (anonymous):

Now, just see that you have + sign with 5 and -ve sign with 1.. So, sign you will not interchange, yes you can interchange the terms but not the sign.. See here: Can I write this (-1) + (5) as: (5) + (-1)

OpenStudy (anonymous):

okay so i can change the format it is, but dont change the sign at all.

OpenStudy (anonymous):

See there, 5 is positive and - sign I have attached to 1.. And please note that if there is no sign before any number, that number is by default positive I mean (5) = +(5) these both are one and the same..

OpenStudy (anonymous):

ok i see that =]

OpenStudy (anonymous):

Yes you can change the format, but not the sign.. Now let us see how to deal with situation where you have two signs: See, you have + (-1) there, right??? (Forget about 5 for sometime)

OpenStudy (anonymous):

Here you have two signs, first + and then - with 1.. Right??

OpenStudy (anonymous):

right

OpenStudy (anonymous):

So, just remember these rules: +(-a) = -a (When + is outside and - is inside, then -will win) -(+a) = -a (here also out of - outside and + inside, - will win) -(-a) = +a (If you have two - signs, then it becomes +) +(+a) = +a (+ and + remain +)

OpenStudy (anonymous):

So, you have +(-1) So now seeing above rules, try to tell me: \(\large + (-1) = ??\)

OpenStudy (anonymous):

-1 right?

OpenStudy (anonymous):

Good...

OpenStudy (anonymous):

And suppose you have : -(-1) In this case what you will get?? (Just look above rules)

OpenStudy (anonymous):

+1

OpenStudy (anonymous):

So, I assume you are getting it correctly.. Now tell me can I write (5) + (-1) as (5) - 1 ??

OpenStudy (anonymous):

no because the one is neg

OpenStudy (anonymous):

Or Simply as 5 is positive so, there is no need to keep brackets with 5 : \(\large 5 - 1 = ??\)

OpenStudy (anonymous):

See, you have solved the last part right now.. you yourself told me that +(-1) = -1, Right??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then you can write -1 there now.. See I show you clearly, just look for red colored part: \[ \large 5 \color{red}{+ (-1)} \implies 5 \color{red}{-1} = How \; \; much \; \; ??\]

OpenStudy (anonymous):

4

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

How minus??

OpenStudy (anonymous):

5 is greater than 1.. So, sign will be of greater number... Since 5 has positive sign so, sign will also be positive...

OpenStudy (anonymous):

ph ok yes i get it

OpenStudy (anonymous):

oh*

OpenStudy (anonymous):

See, you have 5 chocolates, you ate 1 chocolate, now how many chocolates you have??

OpenStudy (anonymous):

4

OpenStudy (anonymous):

\(5 -1 = 4\) \(1 - 5 = -4\) Getting why?

OpenStudy (anonymous):

sure do =] finally

OpenStudy (anonymous):

Here in 1 - 5, it means you have 1 chocolate, and you ate 5 chocolates, But HOW IT IS POSSIBLE.. So, here also, you will do 5 - 1, but as 5 has -ve sign, so, -ve sign you will give to final answer too... Hope you are getting it..

OpenStudy (anonymous):

Coming back to our question: In numerator: You got : \[\large m^{-1} \cdot m^5 = (m)^{\color{red}{-1 + 5}} = (m)^\color{red}{4}\]

OpenStudy (anonymous):

Getting??

OpenStudy (anonymous):

so in the red would it be me adding the -1+5 and the 4?

OpenStudy (anonymous):

So, now we have: \[\large (\frac{m^4}{m^{-2}})^3\]

OpenStudy (anonymous):

I have added in the very next term, see there I have written 4..

OpenStudy (anonymous):

See, we have solved somewhat, In numerator you have two m terms, now in numerator you have one m only..

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Having difficulty anywhere till now, you must ask me right now...

OpenStudy (anonymous):

no

OpenStudy (anonymous):

So. let us proceed further: Our second formula that we have to apply here is: \[\huge \frac{a^x}{a^y} = a^{x - y}\]

OpenStudy (anonymous):

See, you have one m in the numerator and now one m in the denominator, Right??

OpenStudy (anonymous):

Please note: Earlier, you have two m in multiplication like (m.m) in this case, you have to add powers.. But when the two m will be in division like (m/m) then you will subtract the powers in this case.. Are you getting the difference between these two statements?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So now you have powers as 4 in numerator and (-2) in denominator.. So, you will subtract them like this : Please note the red sign: \[\large 4 \color{red}{-} (-2)\]\ Getting??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

PLEASE NOTE: This you must remember that numerator power will come first, then - sign in between and then denominator power.. Don't ever do this : \(\large (-2) \color{red}{-} (4)\) THIS IS WRONG..

OpenStudy (anonymous):

So please be careful here...

OpenStudy (anonymous):

Suppose you have : \(\huge \frac{m^{-2}}{m^4}\) YES, IN THIS CASE YOU CAN SURELY DO : (-2) - (4) as numerator power is now (-2) and denominator power is 4..

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Now tell me in 4 - (-2) \(- (-2) = ??\)

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Or +2 right??

OpenStudy (anonymous):

right

OpenStudy (anonymous):

So, you can do like this now: Look for the red part here also: \[\large 4 \color{red}{- (-2)} \implies 4 \color{red}{+2} = ??\]

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Yep.. Getting how we did it??

OpenStudy (anonymous):

yeppers =]

OpenStudy (anonymous):

So, are we left with now: \[\large \color{blue}{(m^6)^3}\]

OpenStudy (anonymous):

We have solved somewhat further, we had one m upside and one m downside, and now by solving that also, we have now one m...

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Now third formula which you have to remember here which is simple also is: \[\large \color{green}{(a^x)^y = a^{xy}}\] In this case powers will get simply multiplied..

OpenStudy (anonymous):

So, just try and tell me, can you reach your answer from here??

OpenStudy (anonymous):

You are just one step away to reach there, just try otherwise I am here only..

OpenStudy (anonymous):

Hey, just try and tell the answer, at the most you can be right only, nothing greater than that... :)

OpenStudy (anonymous):

so the last part is to mul

OpenStudy (anonymous):

Yep, go ahead..

OpenStudy (anonymous):

6? i feel that is wrong though....

OpenStudy (anonymous):

x = 6 and y = 3 that is outside, multiply them just..

OpenStudy (anonymous):

18

OpenStudy (anonymous):

Yes you are there.. but 18 is in the power of m...

OpenStudy (anonymous):

so would it be 1/m18?

OpenStudy (anonymous):

List of Fomulae that we have used here: \[\large \color{green}{1. \; \; a^x \cdot a^y = a^{x+y}}\] \[\large \color{red}{2. \; \; \frac{a^x}{a^y} = a^{x-y}}\] \[\large \color{blue}{(a^x)^y = a^{x \cdot y}}\]

OpenStudy (anonymous):

See, it will not go in denominator, as power is 18 and that is positive..

OpenStudy (anonymous):

\(m^{18}\) will remain in numerator.. Yes if you would have -18 in the power, then it would be what you are thinking..

OpenStudy (anonymous):

just m^18?

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

\[\huge \color{green}{(m)^{-a} = \frac{1}{(m)^a}}\]

OpenStudy (anonymous):

okay I get it that is what I was thinking it was at first but i wasnt sure because the problem had the (/) sign and i wasnt sure if i needed to keep it. thats is why i asked if it was 1/m18

OpenStudy (anonymous):

but i think that you have helped me more than any math teacher thank you so much!

OpenStudy (anonymous):

No \ sign we had removed when we did: \[\large \color{green}{ \frac{m^4}{m^{-2}} = m^{4 + 2} = m^6}\]

OpenStudy (anonymous):

See, since you are somewhat new to this topic, I am providing you whole solution step by step, Just go through it and learn how we did it...

OpenStudy (anonymous):

We have our problem: \(\huge (\frac{m^{-1} \cdot m^5}{m^{-2}})^3\) \(\huge \implies (\frac{m^{-1 + 5}}{m^{-2}})^3\) \(\huge \implies (\frac{m^4}{m^{-2}})^3\) \(\huge \implies (m^{4 - (-2)})^3\) \(\huge \implies (m^{4+2})^3\) \(\huge \implies (m^6)^3\) \(\huge \implies (m)^{6 \cdot 3}\) \(\huge \color{blue}{\implies (m)^{18}}\)

OpenStudy (anonymous):

If you have any doubt in any step, just let me know, but the main thing is, you should get each and every step, because these formulae and rules, which I have told you will be used frequently in next problems you will do... So, try more and more problems, I am sure you will get familiar as well as comfortable with these... Don't afraid of these, just make them your friends, they are real easy...

OpenStudy (anonymous):

*be

OpenStudy (anonymous):

can you check to see if i have the right answer on another one i just did on my own?

OpenStudy (anonymous):

I can check only, hurry up because I have to have my dinner now.. Be quick..

OpenStudy (anonymous):